OpenStudy (anonymous):

Log base 7 (3x-11) find domain and write in interval notation

OpenStudy (anonymous):

\[\log_{7}(3x-11) \]

myininaya (myininaya):

can you tell me what the domain is for y=log(x)?

OpenStudy (anonymous):

0? or infinity?

myininaya (myininaya):

it is a range of numbers not just one number

OpenStudy (anonymous):

Then I am lost, sorry

myininaya (myininaya):

how about a graph of log(x)=y recall the domain is the set of x values where the relation (function in this case) exists

myininaya (myininaya):

looking at my graph for y=log(x) can you tell me for what x values the function exists?

OpenStudy (anonymous):

at 3 or 4? cant tell without numbers

OpenStudy (phi):

myininaya is showing that the graph of ln(x) exists only for x>0 ln(0) will give you a math error on your calculator. that is because e^x =0 is not true for any x

myininaya (myininaya):

ok that vertical line represents x=0 look at that graph it continues for much greater x values

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So it will be infinite positive numbers until we reach x?

myininaya (myininaya):

the set is positive numbers or you could say (0,infinity)

myininaya (myininaya):

both of these say things that we can only have positive numbers for our domain anyways we are looking at y=log(f(x)) we need f(x) to be positive which means we need f(x)>0

OpenStudy (anonymous):

so 3x-11>0?

myininaya (myininaya):

omg you took the words right out of my mouth good job

myininaya (myininaya):

solve that for x and you are done

myininaya (myininaya):

well i guess you are also require to put it in interval notation after that

OpenStudy (anonymous):

Lol sometimes words just confuse me. I am better with numbers i feel. so if i can find the proper way to set it i can complete it you know?

OpenStudy (anonymous):

And yes. Can you explain to me how do go about doing so? My first guess would be (11/3,infinity)

myininaya (myininaya):

that is right

OpenStudy (phi):

To review, the domain of a function are the numbers it "accepts" (without causing an error or being undefined). Lots of functions have a domain of all real numbers. Some, like the log, only accept numbers bigger than zero.

OpenStudy (anonymous):

Lol thank u guys, i feel so slow until i actually get going. can I try one more with u to make sure i got it?

myininaya (myininaya):

do you want me to make up one or do you have one?

OpenStudy (anonymous):

I am taking practice problems from the book so i have one.

OpenStudy (anonymous):

\[\log (x-5)+\log(8-.5x)\]

OpenStudy (anonymous):

This might seem very elementary but when there is no base it is recognized as 10 right?

OpenStudy (phi):

yes

myininaya (myininaya):

this is another domain question?

OpenStudy (anonymous):

yea

myininaya (myininaya):

you can go about this two different ways you can find the domain of log(x-5) and the domain of log(8-.5x) and then find the intersection of those sets to find the domain of log(x-5)+log(8-.5x) or you can find condense into one log and then remember that if you have y=log(f(x)) the domain is f(x)>0

myininaya (myininaya):

do you know what i mean by intersection?

OpenStudy (anonymous):

Im pretty sure i do. I think id like the first way a little better.

myininaya (myininaya):

so what is the domain of log(x-5)?

OpenStudy (anonymous):

5

OpenStudy (phi):

the domain of a function are the numbers it "accepts"

OpenStudy (anonymous):

\[-\infty,5\]

myininaya (myininaya):

so can i ask what inequality you solved to get that?

OpenStudy (anonymous):

uh oh lol , means im wrong

OpenStudy (anonymous):

x-5=0

myininaya (myininaya):

is that an inequality?

OpenStudy (anonymous):

Dangit x-5>0?

myininaya (myininaya):

ok great and you added 5 on both sides right do you know how to read x>5?

myininaya (myininaya):

like what does that mean?

OpenStudy (anonymous):

All numbers greater than 5 are allowed?

myininaya (myininaya):

so you mean to say the domain of log(x-5) is (5,inf)?

myininaya (myininaya):

inf=infinity

OpenStudy (anonymous):

Yes, i think lol

myininaya (myininaya):

it is (5,inf) means 5 and any number after 5 without actually including 5

myininaya (myininaya):

so any number after 5 lol